# Prospect of Precision Cosmology and Testing General Relativity using Binary Black Holes- Galaxies Cross-correlation

### v1

@@ 6.4 Reconstruction of F(z) and Hubble Constant
By extending the hierarchical Bayesian framework used in the previous section to measure $\rm{\mathcal{F}(z)}$ with a fixed $\rm{H_{0}}$, we can also jointly measure $\rm{\mathcal{F}(z)}$ and $\rm{H_{0}}$. In our approach, we determine the EM luminosity distance from various length scales, including the BAO scale, sound horizon distance, and redshift. Figures 11 and 12 provide a comprehensive comparison of $\rm{\mathcal{F}(z)}$ measurements under the scenario of varying $\rm{H_{0}}$ for both the LVK and CEET sources respectively by combining with DESI and LSST. This comparative analysis highlights the impact of galaxy surveys on the precision of $\rm{\mathcal{F}(z)}$ measurements. Additionally, it underscores the potential of advanced detectors like CEET to deliver increasingly accurate cosmological insights, extending deep into higher redshifts. We also show in Figure 13 the optimal combined measurements of $\rm{\mathcal{F}(z)}$ and the Hubble constant ($\rm{H_{0}}$) at $\rm{z=0.425}$ for both the CEET and LVK systems with the DESI galaxy survey. It shows that there exists a degeneracy between $\rm{H_{0}}$ and $\rm{\mathcal{F}(z)}$, as a result, the error on the individual parameters increases. Allowing $\rm{H_{0}}$ to vary increases the errors in the measurements of $\rm{\mathcal{F}(z)}$ by a factor of $\sim$ 2 compared to measurements with a fixed $H_{0}$. The typical error on the measurements of $\rm{\mathcal{F}(z)}$ with a varying $\rm{H_{0}}$ scenario for LVK+DESI ranges around 19% between $\rm{z=0.325}$ and $\rm{z=0.625}$, while for LVK+LSST it is around 23%. Outside this range, the error increases significantly, as shown in Figure 11. Similarly, for CEET+DESI, the typical error is around 7.5% between $\rm{z=0.275}$ and $\rm{z=1.2}$, and for CEET+LSST, it is around 9%. Beyond these ranges, the error also increases significantly, as depicted in Figure 12.
@@
The primary source of uncertainty in $\rm{\mathcal{F}(z)}$ measurements stems from redshift errors, which are estimated through the cross-correlation between GW sources and galaxy surveys. The sky coverage of these galaxy surveys is crucial in this context. As the survey size increases, the redshift error decreases, thereby enhancing the precision of $\rm{\mathcal{F}(z)}$ measurements. Furthermore, the precision of estimating $\rm{\mathcal{F}(z)}$ depends on accurate measurements of both the BAO scale and the GW luminosity distance, $\rm{D_{l}^{GW}(z)}$. Improved accuracy in BAO scale measurements directly enhances the precision of $\rm{\mathcal{F}(z)}$. On the GW side, accurate measurements of $\rm{\mathcal{F}(z)}$ rely on $\rm{D_{l}^{GW}(z)}$ and sky localization. Enhancing the precision of $\rm{D_{l}^{GW}(z)}$ and sky localization consequently improves the accuracy of $\rm{\mathcal{F}(z)}$, as discussed in Subsection 6.2.
@@
In our analysis of measuring the frictional term $\rm{\mathcal{F}(z)}$ and the Hubble constant $\rm{H_{0}}$, several significant sources of uncertainty exist, with one of the most notable being the uncertainty associated with inferring redshifts from the cross-correlation of galaxy surveys with GW sources. Photometric redshift estimation errors from the LSST catalog, due to limited wavelength coverage and color-redshift degeneracies, introduce significant inaccuracies. Although the DESI catalog provides more precise spectroscopic redshifts, minor measurement errors still impact the overall precision. A crucial factor in these measurements is the partial sky coverage of both surveys. The LSST survey covers approximately 18,000 square degrees ($\rm{f_{\text{sky}}\approx 0.436}$), while the DESI survey covers around 14,000 square degrees ($\rm{f_{\text{sky}}\approx 0.339}$). Increasing the sky coverage to a full-sky survey would improve the precision of these measurements by reducing the redshift error by a factor of $\rm{1/\sqrt{f_{\text{sky}}}}$. Combining different surveys that cover varying sky areas can further enhance measurement precision by increasing the effective sky coverage and the number of sources available for cross-correlation. This synergistic approach helps mitigate the limitations of individual surveys and enhances the overall accuracy of cosmological parameter estimations. Additionally, the luminosity distance measurements of GW sources are subject to uncertainties from waveform modeling, instrumental noise, inclination angle degeneracies, and weak lensing. On the other hand, BAO measurements, which serve as a standard ruler for cosmological distances, also contribute to uncertainties due to finite sample sizes and survey geometry. The statistical noise in the cross-correlation process, stemming from the finite number and spatial distribution of galaxies and GW sources, further affects the inferred redshift accuracy.

### v2

@@ 6.4 Reconstruction of F(z) and Hubble Constant
In our approach, we determine the EM luminosity distance from various length scales, including the BAO scale, sound horizon distance, and redshift. Figures 11 and 12 provide a comprehensive comparison of $\rm{\mathcal{F}(z)}$ measurements under the scenario of varying $\rm{H_{0}}$ for both the LVK and CEET sources respectively by combining with DESI and LSST. This comparative analysis highlights the impact of galaxy surveys on the precision of $\rm{\mathcal{F}(z)}$ measurements. Additionally, it underscores the potential of advanced detectors like CEET to deliver increasingly accurate cosmological insights, extending deep into higher redshifts. We also show in Figure 13 the optimal combined measurements of $\rm{\mathcal{F}(z)}$ and the Hubble constant ($\rm{H_{0}}$) at $\rm{z=0.425}$ for both the CEET and LVK systems with the DESI galaxy survey. It shows that there exists a degeneracy between $\rm{H_{0}}$ and $\rm{\mathcal{F}(z)}$, as a result, the error on the individual parameters increases. Allowing $\rm{H_{0}}$ to vary increases the errors in the measurements of $\rm{\mathcal{F}(z)}$ by a factor of $\sim$ 2 compared to measurements with a fixed $H_{0}$. The typical error on the measurements of $\rm{\mathcal{F}(z)}$ with a varying $\rm{H_{0}}$ scenario for LVK+DESI ranges around 19% between $\rm{z=0.325}$ and $\rm{z=0.625}$, while for LVK+LSST it is around 23%. Outside this range, the error increases significantly, as shown in Figure 11. Similarly, for CEET+DESI, the typical error is around 7.5% between $\rm{z=0.275}$ and $\rm{z=1.2}$, and for CEET+LSST, it is around 9%. Beyond these ranges, the error also increases significantly, as depicted in Figure 12.
@@
The primary source of uncertainty in $\rm{\mathcal{F}(z)}$ measurements arises from redshift errors, which are estimated via cross-correlation between GW sources and galaxy surveys. This uncertainty decreases with increased sky coverage of the surveys, improving the precision of $\rm{\mathcal{F}(z)}$ measurements. Accurate estimation of $\rm{\mathcal{F}(z)}$ also depends on precise measurements of both the BAO scale and the GW luminosity distance, $\rm{D_{l}^{\rm{GW}}(z)}$, as well as sky localization. LSST’s photometric redshift errors, due to limited wavelength coverage and color-redshift degeneracies, and DESI’s more precise spectroscopic redshifts, though still affected by minor errors, influence overall precision.Sky coverage is critical; LSST covers approximately 18,000 square degrees ($\rm{f_{\text{sky}}\approx 0.436}$) and DESI around 14,000 square degrees ($\rm{f_{\text{sky}}\approx 0.339}$). Expanding to full-sky coverage would significantly reduce redshift errors. Combining surveys with different sky areas enhances measurement precision by increasing effective sky coverage and the number of sources for cross-correlation, thereby improving cosmological parameter estimation. Additionally, GW luminosity distance measurements face uncertainties from waveform modeling, instrumental noise, inclination angle degeneracies, and weak lensing, while BAO measurements are affected by finite sample sizes and survey geometry. Statistical noise from the limited number and distribution of galaxies and GW sources further impacts inferred redshift accuracy.